OpenStudy (princeevee):

solve the open equation -3<(or equal to)n+6<(or equal to)7 n>(or equal to) -9 and n<(or equal to)1 n<(or equal to)3 and n<(or equal to)7 N>(or equal to)9 and n<(or equal to)7 n<(or equal to) -9 and n<(or equal to)6

OpenStudy (princeevee):



OpenStudy (princeevee):

@Hero

OpenStudy (texaschic101):

-3 <= n + 6 <= 7 --- subtract 6 from ...well...everywhere -3 - 6 <= n + 6 - 6 <= 7 - 6 --- simplify -9 <= n <= 1....so n >= -9 and n <= 1

OpenStudy (texaschic101):

its the 2nd one

OpenStudy (texaschic101):

or you could have split it into 2 inequalities and solved each one.. -3 <= n + 6 and n + 6 <= 7

OpenStudy (princeevee):

k help me with more?

OpenStudy (princeevee):

@Hero

OpenStudy (princeevee):

@Hero